Standardizing NaOH with KHP (molar mass 204.2 g/mole)

mass of flask + KHP 99.829 g final buret reading 25.35 mL

mass of flask 98.326 g initial buret reading 0.05 mL



Mass of KHP ___________ g Volume of NaOH, mL ____________



moles of KHP _____________ moles of NaOH needed neutralize KHP _______________



Molarity of NaOH? _______________
1
Expert's answer
2020-12-03T14:08:23-0500

Standardizing NaOH with KHP (molar mass 204.2 g/mole)


mass of flask + KHP 99.829 g final buret reading 25.35 mL


mass of flask 98.326 g initial buret reading 0.05 mL


204.2g- 99.829g = 104.371g


98.326g- 0.05g = 98.276g of NaOH


1 mole of NaOH = 40g


Therefore :


1mole of NaOH = 40g


? = 98.276g


98.276g/40g = 2.4569moles


= 2.4569moles